40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1292     Entry time: Wed Feb 11 10:52:22 2009
Author: Yoichi 
Type: Configuration 
Category: DAQ 
Subject: C1:PEM-OSA_APTEMP and C1:PEM-OSA_SPTEMP disconnected 

During the cleanup of the lab. Steve found a box with two BNCs going to the ICS DAQ interface and an unconnected D-SUB on the floor under the AP table.  It seemed like a temperature sensor.

The BNCs were connected to C1:PEM-OSA_APTEMP and C1:PEM-OSA_SPTEMP.

Steve removed the box from the floor. These channels can be now used as spare DAQ channels. I labeled those cables.

ELOG V3.1.3-